### v1

@@ 1. Introduction
Cannon–Thurston maps, named after the influential work of Cannon and Thurston in [CT07], are maps induced by the inclusion of a (typically infinite index normal) subgroup $H$ of a group $G$ at the level of boundaries $\partial H\to\partial G$. The most well-studied case is where $H$ and $G$ are hyperbolic, and the boundaries here are the Gromov boundaries. In particular Mitra (Mj) [Mit98], proved that Cannon–Thurston maps exist for any normal hyperbolic subgroup of a hyperbolic group. Cannon–Thurston maps have also been considered in the case of relatively hyperbolic groups and their Bowditch boundary [Pal10, Bow07, Mj14], as well as CAT(0) groups and their visual boundaries [BCG+22].
@@ Acknowledgements
The authors would like to thank SNSF Ambizione Fellowship of the second author used to organize a research retreat where this paper had its genesis. The second author was supported in part by an SNSF Ambizione Fellowship. The third author was supported by both the EPSRC DTA studentship as well as the LMS Early Career Fellowship.
@@ 2. Preliminaries
For a detailed definition of the Morse boundary, see [Cor17, CH17] and the corrigendum [CSS23].
@@ 3.1. Elements fixed by automorphisms
We start with a preliminary lemma.
Lemma 3.4.
Let $G=H\rtimes_{\varphi}Q$. Let $q\in Q$ be an infinite-order undistorted element of $Q$, and let $h\in H$ be an infinite order element such that $\varphi_{q}(h)=h$. There exist quasi-geodesic constants $(K,C)$ such that the following holds: for every $D>0$ there exists $k_{0}>0$ such that for all $k\geq k_{0}$ there exists a pair $\gamma_{1}$ and $\gamma_{2}$ of $(K,C)$–quasi-geodesics from $1$ to $q^{k}$ such that the Hausdorff distance between $\gamma_{1}$ and $\gamma_{2}$ at least $D$.
Proof.
Note that if $q$ is undistorted in $Q$, then it is also undistorted in $G$, since there is a homomorphism $G\to Q$ restricting to the identity on $Q$.
Observe that $d_{G}(h^{n},\{q^{k}\}_{k\in\mathbb{N}})\to\infty$ for $n\to\infty$. Indeed, if this was not the case, there would exists a constant $D>0$ and a sequence $(k_{n})_{n}\in\mathbb{N}$ such that $d_{G}(h^{n},q^{k_{n}})<D$ for all $n\in\mathbb{N}$. In other words, $d_{G}(1,q^{-k_{n}}h^{n})<D$. Since the Cayley graph is locally finite, there then exists $n\neq m$ such that $h^{n}q^{-k_{n}}=h^{m}q^{-k_{m}}$ implying that $1\neq h^{n-m}=q^{k_{n}-k_{m}}$, a contradiction to the definition of the semi-direct product.
Since $\langle q\rangle$ is undistorted, the continuous map $\gamma:\mathbb{R}\to G$ defined by the bi-infinite concatenation of geodesics
$\ldots[q^{-2},q^{-1}]\circ[q^{-1},1]\circ[1,q]\circ[q,q^{2}]\ldots$
is a $(K^{\prime},C^{\prime})$–quasi-geodesic for some constants $(K^{\prime},C^{\prime})$. We prove that the lemma holds for $(K,C)=(3K^{\prime},C^{\prime})$.
Let $D>0$. Let $D_{1}=2\max\{D+d_{G}(1,q),d_{G}(1,h)\}+1$ and let $D_{2}$ be large enough such that the following works.
Let $n\in\mathbb{N}$ such that $D_{1}/2\leq d_{G}(h^{n},\{q^{k}\}_{k\in\mathbb{Z}})\leq D_{1}$. Such an $n$ exists by the observation above. Note that for any $l\in\mathbb{N}$, we have that $d_{G}(h^{n}q^{l},\gamma)\geq D_{1}/2-d_{G}(1,q)>D$.
Choose $k_{0}\geq 0$ such that $d_{G}(1,q^{k})\geq D_{2}$ for all $k\geq k_{0}$. Let $k\geq k_{0}$ and let $x=h^{n}\gamma(t)$ and $y=h^{n}\gamma(t^{\prime})$ be the points on $h^{n}\gamma$ closest to $1$ and $q^{k}$. By Lemma 2.9 of [Zbi23], the path $\gamma_{1}=[1,x]\circ h^{n}\gamma\mid_{[t,t^{\prime}]}\circ[y,q^{k}]$ is a $(K,C)$–quasi-geodesic for large enough $D_{2}$. Further, as shown above, there exists $l\in\mathbb{Z}$ such that $h^{n}q^{l}\in h^{n}\gamma\mid_{[t,t^{\prime}]}$, implying that the Hausdorff distance between $\gamma_{1}$ and $\gamma_{2}=\gamma|_{[1,q^{k}]}$ is more than $D$. ∎
@@
Let $(K,C)$ be the constants from Lemma 3.4 and let $M$ be a Morse gauge. In light of Lemma [Cor17], there exists a constant $D$ such that any pair $\gamma_{1},\gamma_{2}$ of $(K,C)$–quasi-geodesics with the same endpoints on an $M$–Morse geodesic have Hausdorff distance at most $D$. Lemma 3.4 shows that for large enough $k$, there exist two $(K,C)$–quasi-geodesics from $1$ to $q^{k}$ which are at Hausdorff distance more than $D$. Hence, for large enough $k$, $[1,q^{k}]$ is not $M$–Morse. This implies that for large enough $L$, if $d(q^{l},q^{k+l})\geq L$, then $[q^{l},q^{k+l}]=q^{l}[1,q^{k}]$ is not $M$-Morse, which concludes the proof. ∎
@@ 3.2. Iterates are not weakly Morse
Assume by contradiction that $\{\varphi_{q}^{n}(h)\}_{n\in\mathbb{N}}$ is weakly $M$–Morse for some Morse gauge $M$. In fact, we can assume that any subsegment of $[1,\varphi_{q}^{n}(h)]$ is $M$-Morse (see for example Lemma 2.1 of [Cor17]). Let $C=M(3,0)$ and let $M^{\prime}$ be the Morse gauge such that in every quadrangle where one side has length at most $C$, two of the other sides are $M$-Morse, the third side is $M^{\prime}$–Morse (such an $M^{\prime}$ exists by [Cor17, Lemma 2.3]). Let $L\geq 4d_{G}(1,q)$ be the length such that for every pair of points $q^{k},q^{l}$ with $d_{G}(q^{k},q^{l})\geq L$ we have that $[q^{k},q^{l}]$ is not $M^{\prime}$–Morse. Such an $L$ exists since $q$ is not pairwise Morse. Lastly, let $n\in\mathbb{N}$ be such that $d_{G}(1,\varphi_{q}^{n}(h))\geq 10^{9}L$ and let $k$ be such that $L\leq d_{G}(1,q^{k})\leq 2L$.
@@ 3.3. A criterion for non-existence of Cannon-Thurston maps
Assume by contradiction that there exists a continuous $H$–equivariant map $f:\partial_{*}H\to\partial_{*}G$. Let $a\in\partial_{*}H$. Since $\{h_{n}\}_{n}$ is weakly Morse in $H$, there exists a Morse gauge $M$ such that $h_{n}a\in\partial_{*}^{M}H$ for all $n\in\mathbb{N}$ (this is because if two sides of a triangle are Morse, then the third one is as well, quantitatively, see [Cor17, Lemma 2.3]). We can assume that $M$ is refined in the sense of [CSS23], so that the stratum $\partial_{*}^{M}H$ is compact. The sequence $(h_{n}\cdot a)_{n}$ has a converging subsequence. In fact, up to passing to a subsequence, we can assume that $(h_{n}\cdot a)_{n}$ converges to a point $a_{\infty}\in\partial_{*}H$.
@@ 3.4. An example
First of all, $H$ is hyperbolic relative to $\mathbb{Z}^{2}$, and therefore an element of $H$ is Morse if and only if it has infinite order and is not conjugate into $\mathbb{Z}^{2}$ (this follows from the fact that elementary closures of hyperbolic elements of relatively hyperbolic groups can be added to the list of peripherals [Osi06, Corollary 1.7], combined with strong quasiconvexity of peripheral subgroups [DS05, Lemma 4.15] ). Since $\psi=\varphi|_{H_{0}}$ is hyperbolic, we have that $G_{0}=H\rtimes_{\psi}\mathbb{Z}$ is hyperbolic by [BF92]. Using the decomposition $G=G_{0}\ast_{\langle t\rangle}\mathbb{Z}^{3}$ and the combination theorem for relative hyperbolicity [Dah03], we then see that $G_{0}$ is hyperbolic relative to the mapping torus $\mathbb{Z}^{3}$ of $\mathbb{Z}^{2}<H$. If we have a Morse element $h$ of $H$, in $G$ it is not conjugate into $\mathbb{Z}^{3}$, and therefore it is Morse in $G$ (again by [Osi06, DS05]), concluding the proof. ∎
@@
We can compare this with the case when $H$ is a free group and $\varphi$ is the atoroidal automorphism defined by $a\mapsto aba$ and $b\mapsto ba$. In this case the the Morse Cannon–Thurston map does not exist [BCG+22, Proposition 5.5]. A standard calculation shows that $[a,b]$ commutes with $\langle t\rangle$ in $G=F_{2}\rtimes\langle t\rangle$. So here, the map is not type preserving, and that is where the failure of the Cannon-Thurston map is exhibited. In fact, all examples covered by [BCG+22] fail to be type preserving.
@@ 4. A mapping torus of a RAAG with a Cannon-Thurston map
There exists an equivariant isometry $T\to T^{\prime}$.
@@
We have that $H^{(M)}$ quasi-isometrically embeds into $T$ by [CCS23, Lemma 4.5]. Therefore, in view of Lemma 4.2, we have that the composition of $\iota_{M}$ and an orbit map $\pi$ of $G$ to $T^{\prime}$ is a quasi-isometric embedding. Since $\pi$ is coarsely Lipschitz, $\iota_{M}$ needs to be a quasi-isometric embedding. Stability follows from acylindricity of the action, Lemma 4.3, which ensures that geodesics $\gamma$ in $G$ with the property that $\pi\circ\gamma$ is a quasi-geodesic are Morse (with Morse gauge controlled in terms of the hyperbolicity constant); this is Lemma [CCS23, Lemma 4.6] (see also [DMS10, Sis16]). ∎

### v2

@@ 1. Introduction
Cannon–Thurston maps, named after the influential work of Cannon and Thurston in [CT07], are maps induced by the inclusion of a (traditionally infinite index normal) subgroup $H$ of a group $G$ at the level of boundaries $\partial H\to\partial G$. The most well-studied case is where $H$ and $G$ are hyperbolic, and the boundaries here are the Gromov boundaries. In particular Mitra (Mj) [Mit98], proved that Cannon–Thurston maps exist for any normal hyperbolic subgroup of a hyperbolic group. Cannon–Thurston maps have also been considered in the case of relatively hyperbolic groups and their Bowditch boundary [Pal10, Bow07, Mj14], as well as CAT(0) groups and their visual boundaries [BCG+22].
@@
As we will see in the proof of Theorem 1.2, the reason that the Cannon-Thurston map exists in this case is the interesting phenomenon that, while $H$ is distorted in $G$, all Morse directions of $H$ are in fact undistorted in $G$.
@@ 1.1. Some open questions
Note that JSJ theory puts strong restrictions on $H$ and $G$ as above, but on the other hand, if one such subgroup exists, a purely pseudo-Anosov but not convex-cocompact subgroup $H$ of the mapping class group of a closed surface $\Sigma$ could give rise to an example. Indeed, given $H$ as above there is a short exact sequence $1\to\pi_{1}(\Sigma)\to G\to H\to 1$, and $G$ is not hyperbolic but it does not contain $\mathbb{Z}^{2}$-subgroups, see [FM02, Ham05]. This second fact is not sufficient to show that non-trivial elements of $H$ are Morse in $G$, and in fact we do not know how to prove this, but it can be regarded as evidence towards this.
@@ Acknowledgements
The authors would like to thank SNSF Ambizione Fellowship of the second author used to organize a research retreat where this paper had its genesis. The second author was supported in part by an SNSF Ambizione Fellowship. The third author was supported by both the EPSRC DTA studentship as well as the LMS Early Career Fellowship. All authors would like to thank the referee for prompt and helpful comments.
@@ 2. Preliminaries
For a detailed definition of the Morse boundary, see [Cor17, CH17] and the corrigendum [CSZ24].
@@ 3. Non-existence of Cannon–Thurston maps
In the statement below we use the notation $H\rtimes_{\varphi}Q$ for the semidirect product given by a homomorphism $\varphi:Q\to\textrm{Aut}(H)$. For $q\in Q$, we denote $\varphi_{q}=\varphi(q)$.
@@ 3.1. Elements fixed by automorphisms
Note that if $q$ is an infinite-order undistorted element of $Q$, then if it is not Morse, it is also pairwise not Morse. As shown in e.g. [CW21, Lemma 3.8], if $q$ is Morse, then $\langle q\rangle$ has finite index in its centralizer. If $\varphi_{q}(h)=h$, then $\langle h\rangle$ is in the centralizer of $q$ and hence $q$ cannot be Morse. ∎
@@ 3.2. Iterates are not weakly Morse
Assume by contradiction that $\{\varphi_{q}^{n}(h)\}_{n\in\mathbb{N}}$ is weakly $M$–Morse in $G$ for some Morse gauge $M$. In fact, we can assume that any subsegment of $[1,\varphi_{q}^{n}(h)]$ is $M$-Morse (see for example Lemma 2.1 of [Cor17]). Let $C=M(3,0)$ and let $M^{\prime}$ be the Morse gauge such that in every quadrangle where one side has length at most $C$, two of the other sides are $M$-Morse, the fourth side is $M^{\prime}$–Morse (such an $M^{\prime}$ exists by [Cor17, Lemma 2.3]). Let $L\geq 4d_{G}(1,q)$ be the length such that for every pair of points $q^{k},q^{l}$ with $d_{G}(q^{k},q^{l})\geq L$ we have that $[q^{k},q^{l}]$ is not $M^{\prime}$–Morse. Such an $L$ exists since $q$ is not pairwise Morse. Lastly, let $n\in\mathbb{N}$ be such that $d_{G}(1,\varphi_{q}^{n}(h))\geq 10^{9}L$ and let $k$ be such that $L\leq d_{G}(1,q^{k})\leq 2L$.
@@ 3.3. A criterion for non-existence of Cannon-Thurston maps
Assume by contradiction that there exists a continuous $H$–equivariant map $f:\partial_{*}H\to\partial_{*}G$. Let $a\in\partial_{*}H$. Since $\{h_{n}\}_{n}$ is weakly Morse in $H$, there exists a Morse gauge $M$ such that $h_{n}a\in\partial_{*}^{M}H$ for all $n\in\mathbb{N}$ (this is because if two sides of a triangle are Morse, then the third one is as well, quantitatively, see [Cor17, Lemma 2.3]). We can assume that $M$ is refined in the sense of [CSZ24], so that the stratum $\partial_{*}^{M}H$ is compact. The sequence $(h_{n}\cdot a)_{n}$ has a converging subsequence. In fact, up to passing to a subsequence, we can assume that $(h_{n}\cdot a)_{n}$ converges to a point $a_{\infty}\in\partial_{*}H$.
@@ 3.4. An example
First of all, $H$ is hyperbolic relative to $\mathbb{Z}^{2}$, and therefore we claim:
Claim 1.
An element $h$ of $H$ is Morse if and only if it has infinite order and is not conjugate into $\mathbb{Z}^{2}$.
Proof.
Morse elements in any group have infinite order and cannot be contained in $\mathbb{Z}^{2}$ subgroups, see e.g. [CW21, Lemma 3.8], so we only need to show the “if” part. By [Osi06, Corollary 1.7], we have that $h$ is contained in an elementary subgroup $E(h)$ with the property that $G$ is hyperbolic relative to $\{\mathbb{Z}^{2},E(g)\}$. Therefore, $E(g)$ is a peripheral subgroup of a relatively hyperbolic structure on $H$, and peripheral subgroups are strongly quasiconvex by [DS05, Lemma 4.15]. This readily implies Morseness of $h$ (since $\langle h\rangle$ and $E(h)$ lie at finite Hausdorff distance). ∎
Since $\psi=\varphi|_{H_{0}}$ is hyperbolic, we have that $G_{0}=H\rtimes_{\psi}\mathbb{Z}$ is hyperbolic by [BF92]. Using the decomposition $G=G_{0}\ast_{\langle t\rangle}\mathbb{Z}^{3}$ and the combination theorem for relative hyperbolicity [Dah03], we then see that $G_{0}$ is hyperbolic relative to the mapping torus $\mathbb{Z}^{3}$ of $\mathbb{Z}^{2}<H$. If we have a Morse element $h$ of $H$, in $G$ it is not conjugate into $\mathbb{Z}^{3}$, and therefore it is Morse in $G$ (again by [Osi06, DS05]), concluding the proof. ∎
@@
We can compare this with the case when $H$ is a free group and $\varphi$ is the atoroidal automorphism defined by $a\mapsto aba$ and $b\mapsto ba$. In this case the Morse Cannon–Thurston map does not exist [BCG+22, Proposition 5.5]. A standard calculation shows that $[a,b]$ commutes with $\langle t\rangle$ in $G=F_{2}\rtimes\langle t\rangle$. So here, the map is not type preserving, and that is where the failure of the Cannon-Thurston map is exhibited. In fact, all examples covered by [BCG+22] fail to be type preserving.
@@ 4. A mapping torus of a RAAG with a Cannon-Thurston map
There exists a $G$-equivariant isometry $T\to T^{\prime}$.
@@
We have that $H^{(M)}$ quasi-isometrically embeds into $T$ by [CCS23, Lemma 4.5]. Therefore, in view of Lemma 4.2, we have that the composition of $\iota_{M}$ and an orbit map $\pi$ of $G$ to $T^{\prime}$ is a quasi-isometric embedding. Since $\pi$ is coarsely Lipschitz, $\iota_{M}$ needs to be a quasi-isometric embedding. Stability follows from acylindricity of the action, Lemma 4.3, which ensures that geodesics $\gamma$ in $G$ with the property that $\pi\circ\gamma$ is a quasi-geodesic are Morse (with Morse gauge controlled in terms of the hyperbolicity constant); this is[CCS23, Lemma 4.6] (see also [DMS10, Sis16]). ∎